# Weighted Point Configurations with Hyperuniformity: An Ecological Example and Models

### v1

@@ 4.2 Results of numerical simulations
Figure 8: Dependence on the number of iterations $T$ is shown in the $\log R_{\ell_{x}}$ versus $\log\ell_{x}$ plots, where $L=L_{x}=L_{y}=4000$ and $a_{0}\equiv\theta_{0}=1000$ for (4.18).
Based on the observations in bush 1–5 reported in Section 3.5 such that the distribution of mass $M$ is well approximated by the exponential distribution, first we choose the exponential distribution for $A$, since the mass transform $M$ in each coalescing (4.17) will be given by the area $A$ multiplied by density of points. The probability density function in (4.15) is assume to be given by
$p(a)=p(a;\theta_{0})=\frac{1}{\theta_{0}}e^{-a/\theta_{0}},$ (4.18)
where $\theta_{0}$ is the scale parameter. It is easy to verify that $\langle A\rangle=\theta_{0}\equiv a_{0}$ and $\mathrm{Var}[A]=\theta_{0}^{2}$. For simplicity, in the computer simulations of our model, we fix the aspect ratio $\lambda=1$; that is, $L_{x}=L_{y}\equiv L$.
Let $L=4000$ and $a_{0}=1000$. Starting from a sample of PPP obtained in Section 3.2, we have iterated the transformation $\mathcal{T}$. For each number of iterations $T$, we have performed the measurements explained in Section 3.3. The $\log R_{\ell_{x}}$ versus $\log\ell_{x}$ plots are given in Fig. 8 from $T=0$ (the initial PPP) to $T=500$. As $T$ increases, the value of $\log R_{\ell_{x}}$ increases and the region of $\log\ell_{x}$ in which $\log R_{\ell_{x}}$ shows a linear decay seems to extend. Moreover, we have observed convergence of the plots when $T\gtrsim 500$. We also confirmed the convergence of the plots to the same curve in the simulations starting from a sample of GPP prepared in Section 3.2. From now on, we will report the results of numerical simulations, which were obtained after $T=500$ iterations of the transformation $\mathcal{T}$.
@@ 4.2.1 Hyperuniformity
(a) The results for the marked point processes. As $L$ and $a_{0}$ increase, the linear region extends systematically.
(b) The results for the unmarked point processes. Even if the values of $L$ and $a_{0}$ become large, the linear region seems to be restricted in a narrow region of scale.
Figure 9: $\log R_{\ell_{x}}$ versus $\log\ell_{x}$ of the point processes obtained by numerical simulation of the model with variety of system sizes $L=L_{x}=L_{y}$ and $a_{0}$.
@@
Both of the marked point processes $\Pi$ and the corresponding unmarked point processes $\Xi_{\Pi}$ obtained by our computer simulations are systematically studied by changing the system size $L$ and the mean value $a_{0}$ of the random area $A$. Figure 9 (a) shows the $\log R^{\rm mass}_{\ell_{x}}$ versus $\log\ell_{x}$ plots averaged over 20 marked point processes obtained by our model. For fixed $L=2000$, as $a_{0}$ increases from 500 to 4000, the linear region extends systematically. For the largest value $a_{0}=4000$, the linear region also systematically extends as $L$ increases from 1000 to 4000. By these results, we think that if both of $L$ and $a_{0}$ are sufficiently large, and if the number of iterations $T$ of the transformation $\mathcal{T}$ is large enough, our procedure produced marked point processes with hyperuniformity. The linear fitting (3.12) to the plots for $L=4000$ and $a_{0}=4000$ gives
$R_{\ell_{x}}^{\rm mass}\simeq\ell_{x}^{-\alpha^{\rm model}_{\rm mass}}\quad\mbox{with}\quad\alpha^{\rm model}_{\rm mass}\fallingdotseq 0.96.$ (4.19)
Figure 9 (b) shows the $\log R^{\rm point}_{\ell_{x}}$ versus $\log\ell_{x}$ plots averaged over 20 unmarked point processes, which are obtained from the marked point processes by deleting the data of mass distributions. Compared with $\log R^{\rm mass}_{\ell_{x}}$, even if the values of $L$ and $a_{0}$ become large, the linear region seems to be restricted in a narrow region of scale. The linear fitting (3.12) to the plots for $L=4000$ and $a_{0}=4000$ gives
$R_{\ell_{x}}^{\rm point}\simeq\ell_{x}^{-\alpha^{\rm model}_{\rm point}}\quad\mbox{with}\quad\alpha^{\rm model}_{\rm point}\fallingdotseq 0.27.$ (4.20)
We have studied dependence of the results on the choice of distribution (4.15) of random area $A$. As extensions of the exponential distribution (4.18), we consider the cases in which the area $A$ follows the Gamma distributions with the probability density
$p(a)=p(a;k_{0},\theta_{0})=\frac{1}{\Gamma(k_{0})\theta_{0}^{k_{0}}}a^{k_{0}-1}e^{-a/\theta_{0}},$ (4.21)
@@
$\alpha^{\rm model}_{\rm mass}\fallingdotseq 1,\quad\alpha^{\rm model}_{\rm point}\fallingdotseq 0.4.$ (4.22)
Remark 2   As a simplified model, we have simulated the algorithm with a fixed value of area $A$ of disk. Then we found that the number of iterations $T$ needed for convergence of point processes became much larger than 500. In addition, the convergence of the $\log R_{\ell_{x}}$$\log\ell_{x}$ plots to a line needed larger values of $L$ and $a_{0}$ compared to the results shown by Fig. 9. Hence we have concluded that the randomness of area $A$ of disk is important to generate marked point processes with hyperuniformity efficiently.
@@ 4.2.2 Mass distributions
For Figs. 9 (a) and (b), we have performed 20 sets of the $T=500$ iterations of transformation $\mathcal{T}$ starting from independently obtained 20 samples of PPPs. We found that the mass distributions of 20 samples seem to be quite different from each other, especially when $a_{0}$ becomes large. Figure 11 shows two samples of mass distributions obtained by numerical simulations with $L=2000$ and $a_{0}=4000$. The mass distributions depend heavily on samples and they do not seem to be universal.
@@
$p_{\rm G}(m)$ $=p_{\rm G}(m;k_{\rm G},\theta_{\rm G})$
$=\frac{1}{\Gamma(k_{\rm G})\theta_{\rm G}^{k_{\rm G}}}m^{k_{\rm G}-1}e^{-m/\theta_{\rm G}},\quad m\geq 0.$ (4.23)
Here we have fixed the parameters of (4.21) for the model as $a_{0}=4000$ and $L=2000$. The $k_{0}$-dependence of the fitting curves of (4.23) are shown in Fig.12, where fitting parameters $(k_{\rm G},\theta_{\rm G})$ are given by $(2.15,39.29)$, $(2.2,24.26)$, $(2.5,21.6)$, $(2.5,18.13)$, $(2.55,14.48)$, $(2.6,13.0)$, and $(2.6,12.0)$ for $k_{0}=0.5,1,1.5,2,3,4$, and 5, respectively.
Figure 12: The mass distributions averaged over the 20 samples are calculated for several values of $k_{0}$ of the distribution (4.21). Here $L=2000$ and $a_{0}=k_{0}\theta_{0}=4000$. They are well-approximated by (4.23) with appropriate values of $(k_{\rm G},\theta_{\rm G})$.
@@ 5 Discussions and Concluding Remarks
We have succeeded to generate marked point processes with hyperuniformity from uncorrelated PPPs by iterating random thinning-coalescing processes. We have no evidence that similar types of thinning and coalescing processes have been iterated in continuous survival competitions of buses in real deserts. But our numerical simulations suggest that in order to realize hyperuniformity in large scale, sufficiently large number of iterations of local processes should be needed. As shown by Fig. 8, if the number of iteration is not sufficiently large, evident hyerperuniformity will not be achieved. We think that another reason of the small values of $\alpha^{\rm Kenya}=0.36$ and $\alpha^{\rm Algeria}\fallingdotseq 0$ will come from some historical backgrounds of these deserts.
@@
Acknowledgements   AE and MK would like to thank Jun-ichi Wakita for instructing them the image processing techniques to produce digital data of bushes in deserts. AE is grateful to Hiraku Nishimori, Mitsugu Matsushita, Osamu Moriyama, Zouhaïr Mouayn, Helmut R. Brand for useful discussions and encouragement of her researches. MS was supported by JSPS KAKENHI Grant Numbers JP19K03674, JP21H04432, JP22H05105, JP23K25774, and JP24K06888. TS was supported by JSPS KAKENHI Grant Numbers JP20K20884, JP21H04432, JP22H05105, JP23K25774, and JP24KK0060.

### v2

@@ 4.2 Results of numerical simulations
Figure 8: Dependence on the number of iteration-terms $T$ is shown in the $\log R_{\ell_{x}}$ versus $\log\ell_{x}$ plots, where $L=L_{x}=L_{y}=4000$ and $a_{0}=\theta_{0}=1000$ for (18).
Based on the observations in bush 1–5 reported in Section 3.5 such that the distribution of mass $M$ is well approximated by the exponential distribution, first we choose the exponential distribution for $A$, since the mass transform $M$ in each coalescing (17) will be given by the area $A$ multiplied by density of points. The probability density function in (15) is assumed to be given by
$p(a)=p(a;\theta_{0})=\frac{1}{\theta_{0}}e^{-a/\theta_{0}},$ (18)
where $\theta_{0}$ is the scale parameter. It is easy to verify that $a_{0}\equiv\langle A\rangle=\theta_{0}$ and $\mathrm{Var}[A]=\theta_{0}^{2}$. For simplicity, in the computer simulations of our model, we fix the aspect ratio $\lambda=1$; that is, $L_{x}=L_{y}\equiv L$.
Let $L=4000$ and $a_{0}=1000$. We start our algorithm from a sample of PPP obtained in Section 3.2, where $M(\bm{X})=1$ for all points $\bm{X}\in\mathcal{P}$ at the beginning $T=0$. For each marked point process obtained after $T$ iteration-terms of transformation $\mathcal{T}$, we have performed the measurements explained in Section 3.3. The $\log R_{\ell_{x}}$ versus $\log\ell_{x}$ plots are given in Fig. 8 from $T=0$ (the initial PPP) to $T=500$. As $T$ increases, the value of $\log R_{\ell_{x}}$ increases and the region of $\log\ell_{x}$ in which $\log R_{\ell_{x}}$ shows a linear decay seems to extend. Moreover, we have observed convergence of the plots when $T\gtrsim 500$. We also confirmed the convergence of the plots to the same curve in the simulations starting from a sample of GPP with $M(\bm{X})\equiv 1$ prepared in Section 3.2. From now on, we will report the results of numerical simulations, which were obtained after $T=500$ iteration-terms of transformation $\mathcal{T}$.
@@ 4.2.1 Hyperuniformity
Figure 9: $\log R_{\ell_{x}}$ versus $\log\ell_{x}$ of the point processes obtained by numerical simulation of the model with variety of system sizes $L=L_{x}=L_{y}$ and $a_{0}$. (a) For the marked point processes. As $L$ and $a_{0}$ increase, the linear region extends systematically. The linear fitting to (12) for $L=4000$, $a_{0}=4000$ is shown by a black line, whose slope gives $-\alpha^{\rm model}_{\rm mass}\fallingdotseq-0.96$. (b) For the unmarked point processes. Even if the values of $L$ and $a_{0}$ become large, the linear region seems to be restricted in a narrow region of scale. The linear fitting to (12) in the narrow region for $L=4000$, $a_{0}=4000$ is shown by a black line, whose slope gives $-\alpha^{\rm model}_{\rm point}\fallingdotseq-0.27$.
@@
Both of the marked point processes $\Pi$ and the corresponding unmarked point processes $\Xi_{\Pi}$ obtained by our computer simulations are systematically studied by changing the system size $L$ and the mean value $a_{0}$ of the random area $A$. Figure 9 (a) (resp. (b)) shows the $\log R^{\rm mass}_{\ell_{x}}$ (resp. $\log R^{\rm point}_{\ell_{x}}$) versus $\log\ell_{x}$ plots averaged over 20 marked point processes obtained by our model. For $500\leq a_{0}\leq 3000$, $L=2000$, the dots were shifted in the $X\equiv\log\ell_{x}$ direction by $\log\sqrt{4000/a_{0}}$, which corresponds to the dilation of the region $\Lambda_{L}$ so that $a_{0}$ is normalized to be 4000. In (a) the dots with $500\leq a_{0}\leq 3000$, $L=2000$ were shifted also in the $Y\equiv\log R_{\ell_{x}}$ direction appropriately in order to clarify the convergence of slopes of lines along which dots line up.
First we explain the results for the marked point processes shown by Fig. 9 (a). For fixed $L=2000$, as $a_{0}$ increases from 500 to 4000, the linear region extends systematically. For the largest value $a_{0}=4000$, the linear region also systematically extends as $L$ increases from 1000 to 4000. By these results, we think that if both of $L$ and $a_{0}$ are sufficiently large, and if the number of iteration-terms $T$ of the transformation $\mathcal{T}$ is large enough, our procedure produced marked point processes with hyperuniformity. The linear fitting to (12) of the dots of $L=4000$ and $a_{0}=4000$ gives
$R_{\ell_{x}}^{\rm mass}\simeq\ell_{x}^{-\alpha^{\rm model}_{\rm mass}}\quad\mbox{with}\quad\alpha^{\rm model}_{\rm mass}\fallingdotseq 0.96.$ (19)
Next we explain the results for the unmarked point processes. Figure 9 (b) shows the $\log R^{\rm point}_{\ell_{x}}$ versus $\log\ell_{x}$ plots are averaged over 20 unmarked point processes, which are obtained from the marked point processes by deleting the data of mass distributions. Compared with $\log R^{\rm mass}_{\ell_{x}}$, even if the values of $L$ and $a_{0}$ become large, the linear region seems to be restricted in a narrow region of scale. The linear fitting (12) to the plots for $L=4000$ and $a_{0}=4000$ gives
$R_{\ell_{x}}^{\rm point}\simeq\ell_{x}^{-\alpha^{\rm model}_{\rm point}}\quad\mbox{with}\quad\alpha^{\rm model}_{\rm point}\fallingdotseq 0.27.$ (20)
We have studied dependence of the results on the choice of distribution (15) of random area $A$. As extensions of the exponential distribution (18), we consider the cases in which the area $A$ follows the Gamma distributions with the probability density
$p(a)=p(a;k_{0},\theta_{0})=\frac{1}{\Gamma(k_{0})\theta_{0}^{k_{0}}}a^{k_{0}-1}e^{-a/\theta_{0}},$ (21)
@@
$\alpha^{\rm model}_{\rm mass}\fallingdotseq 1,\quad\alpha^{\rm model}_{\rm point}\fallingdotseq 0.4.$ (22)
Remark 2   As a simplified model, we have simulated the algorithm with a fixed value of area $A$ of disk. Then we found that the number of iteration-terms $T$ needed for convergence of point processes became much larger than 500. In addition, the convergence of the $\log R_{\ell_{x}}$$\log\ell_{x}$ plots to a line needed larger values of $L$ and $a_{0}$ compared to the results shown by Fig. 9. Hence we have concluded that the randomness of area $A$ of disk is important to generate marked point processes with hyperuniformity efficiently.
@@ 4.2.2 Mass distributions
For Figs. 9 (a) and (b), we have used 20 samples obtained after $T=500$ iteration-terms of transformation $\mathcal{T}$ starting from independently prepared 20 samples of PPPs. We found that the mass distributions of 20 samples seem to be quite different from each other, especially when $a_{0}$ becomes large. Figure 11 shows two samples of mass distributions obtained by numerical simulations with $L=2000$ and $a_{0}=4000$. The mass distributions depend heavily on samples and they do not seem to be universal.
@@
$p_{\rm G}(\mathsf{m})$ $=p_{\rm G}(\mathsf{m};k_{\rm G},\theta_{\rm G})$
$=\frac{1}{\Gamma(k_{\rm G})\theta_{\rm G}^{k_{\rm G}}}\mathsf{m}^{k_{\rm G}-1}e^{-\mathsf{m}/\theta_{\rm G}},\quad\mathsf{m}\geq 0.$ (23)
Here we have fixed the parameters of (21) for the model as $a_{0}=k_{0}\theta_{0}=4000$ and $L=2000$. The $k_{0}$-dependence of the fitting curves of (23) are shown in Fig.12, where fitting parameters $(k_{\rm G},\theta_{\rm G})$ are given by $(2.15,39.3)$, $(2.20,24.3)$, $(2.50,21.6)$, $(2.50,18.1)$, $(2.55,14.5)$, $(2.60,13.0)$, and $(2.60,12.0)$ for $k_{0}=0.5,1,1.5,2,3,4$, and 5, respectively.
Figure 12: The mass distributions averaged over the 20 samples are calculated for several values of $k_{0}$ of the distribution (21). Here $L=2000$ and $a_{0}=k_{0}\theta_{0}=4000$. They are well-approximated by (23) with appropriate values of $(k_{\rm G},\theta_{\rm G})$.
@@ 5 Discussions and Concluding Remarks
We have succeeded to generate marked point processes with hyperuniformity from uncorrelated PPPs by iterating random thinning-coalescing processes. We have no evidence that similar types of thinning and coalescing processes have been iterated in continuous survival competitions of bushes in real deserts. But our numerical simulations suggest that in order to realize hyperuniformity in large scale, sufficiently large number of iterations of local processes should be needed. As shown by Fig. 8, if the number of iteration is not sufficiently large, evident hyerperuniformity will not be achieved. We think that another reason of the small values of $\alpha^{\rm Kenya}=0.36$ and $\alpha^{\rm Algeria}\fallingdotseq 0$ will come from some historical backgrounds of these deserts.
@@
Acknowledgements   AE and MK would like to thank Jun-ichi Wakita for instructing them the image processing techniques to produce digital data of bushes in deserts. AE is grateful to Hiraku Nishimori, Mitsugu Matsushita, Osamu Moriyama, Zouhaïr Mouayn, Helmut R. Brand for useful discussions and encouragement of her researches. MK was supported by JSPS KAKENHI Grant Numbers JP19K03674, JP21H04432, JP22H05105, JP23K25774, and JP24K06888. TS was supported by JSPS KAKENHI Grant Numbers JP20K20884, JP21H04432, JP22H05105, JP23K25774, and JP24KK0060.